++++ INAF LBC+LUCIFER queue observing, 2011 March 27 - 28

Observer: R. Speziali, L. Pentericci, K. Boutsia
Telescope Operator: G. Becchetti
Support Astronomer: O. Kuhn
Instrument: LBC - LUCIFER

Summary

No clouds or cirrus at the beginning of the night that's why we decide to begin with LBC.
Some problems with mirror equilibrium on the blue side. The first images are elongated.
Seeing highly variable throughout the night going from 0.9 to 2.5''
According to the allseeing script, executed at the end of the night 1/3 of the Buttiglione program could be considered within specs
The Lucifer images should have seeing <1.3'', thus the OBs executed are considered within specs

Data Summary

Project Objects Filter ExpTime Notes
Buttiglione HATLAS B-bessel, z-sloan 1.5h seeing very variable - some elongation problems at the beginning - probably out of specs
Gallerani SDSSJ1048+4637_ref 210zJHK+Z 2hr seeing measure on the image <1.2'' - Tell +ARC+FLAT
Maiolino HDF_K1 200H+K+ OrderSep 1hr Tell +ARC+FLAT
PHOT_STD SA098 UsBg_rz    

Details

Begin night with LBC

12:00 - Bias with the camera not connected to LBT

01:48 - Skyflat
  • Uspec - z : 2sec - 0.5sec PA=0
  • Uspec -z: 3.5sec - 1sec PA=180
  • Bbess - z : 4.4sec - 4sec PA=180
  • Bbess - z : 8.8sec - 8sec PA=0

02:20 - Slew to pointing star WT10_197
  • we have tried to do dofpia before correcting pointing but there is too much coma in Red and spherical in Blue and we cannot collimate
  • Pointing correction OK + co-pointing OK
  • actually we used the new IRAF script lbcrangebal

02:58 - slew to PHOT_STD - SA098 dofpia,
  • it took like 6-8 iterations but it managed to collimate on its own
  • seeing in the blue channel is high and it does not give spherical corrections
  • observed STD in Uspec, B-bessel, g-sloan, r-sloan, z-sloan - airmass 1.25
  • seeing from dimm 1.2''

03:13 - slew to Buttiglione_HATLAS_1 -dofpia
  • Red collimate after 2 iterations, Blue after 3
  • seeing on the blue side still >2.0'' and no spherical corrections were send

03:21 - observing target Buttiglione_HATLAS_1
  • dimm ~1.4'' at the beginning of the OB
  • high ellipticity on the red side
  • all objects seem to be elongated in the same direction
  • after the 4th dither we abort OB and run again dofpia

03:43 - observing again Buttiglione_HATLAS_1
  • now dimm shows >2.0''
  • mirror is not in equilibriun and the blue channel images seem out of focus and comatic

04:26 - dofpia - 2 iterations to converge
  • seeing: Red-1.25 , Blue -1.55

04:30 - observing target Buttiglione_HATLAS_2
  • on the red side the seeing is ~1.3
  • actually the seeing is extremely VARIABLE

05:12 - dofpia - 3 iterations to converge

05:17 - observing target Buttiglione_HATLAS_3
  • seeing is very unstable
  • the second half of the OB seems to have better images
  • we decide to repeat the HATLAS_2, because now maybe the mirror is in equilibrium or the seeing got more stable...

06:03 - dofpia - 1 iteration to converge

06:04 - observing target Buttiglione_HATLAS_2 - repeat
  • the seeing of the first images turned out to be >1.5'' and we decide to abort since it is out of specs

06:10 SWITCH TO LUCIFER

06:46 - Slew to Gallerani SDSSJ1048+4637_ref
  • couldn't find star in image
  • slew to ACT0350 for pointing correction 0001-0002

07:00 - Slew back to Gallerani for acq 0003-0008
  • SG9_spec_acq_science.txt
  • we changed the exp time in 6x10 and we can see the target
  • we have tried to center the target - although the pivot star seems not well centered in the slit

07:19 - SG9_spec_science_Z_1.txt 0009-0020
  • airmass 1.1 at the beginning of the exposure
  • seeing measured on the images ranges from 0.9-1.2''
  • some persistence in the first exposure around the pivot star...

08:27 - SG9_spec_science_Z_2.txt 0021-0032
  • made an Y-9.5'' offset to position object at the initial position
  • seeing measured on the image <1.25''

09:34 - Calib Unit IN
  • ARC 0033-0036
  • FLAT 0037-0048

09:47 - Tell acq for Gallerani 0049-0052

09:58 - Telluric science script for Gallerani 0053-0055

10:02 - MOS acquisition for Maiolino HDF_K1 0056-0059

10:28 - Science script for Maiolino HDF_K1_1 0060-0071
  • seeing in the dimm is very variable from 0.9' to 2.0''

11:35 - Tell acq for Maiolino 0072-0078

11:53 - Telluric science script for Maiolino HDF_K1_1 0079-0084

11:59 - Calib Unit IN - it gave an error initially but it kept moving
  • ARC 0085-0088
  • FLAT 0089-0094

12:35 - EndofNight Script - DoDarks
  • There is noise at the o2dcr darks... but Dave T. said it is of very low level and we should not be worried...

-- KonstantinaBoutsia - 28 Mar 2011
Topic revision: r12 - 03 Apr 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback